find no. of negative integral solution of equation x+y+z= -12

Grade:11

1 Answers

Arun
3 years ago
Dear student
 
Total number of negative integral solution = 
 = C(12-1, 3-1) = C(11,2) = 11 *10/2 = 55
 
Hope you understand.
 
Thanks and regards